Zimmerland has following data is known

Nominal interest rates in 2015 = 2.5 %

CP|2015 = 165

CP|2014 = 150

What is the real interest rate in 2015?


1
Expert's answer
2022-02-22T07:34:42-0500

The inflation rate is:

pi = (165 - 150)/150×100% = 10%.

The real interest rate in 2015 is:

r = i - pi = 2.5 - 10 = -7.5%.